OpenStudy (anonymous):

Trying to figure out the average velocity. Teacher walks 10m North, 3m East. and 6m South in 10 seconds. I got 0.7 m/s. Need Help!!!

OpenStudy (agent0smith):



OpenStudy (agent0smith):

 Can you find x? Hint: the 10m side is equal to 6m + x Then use pythagoras to find d.

OpenStudy (anonymous):

x=4m d= 3^2+4^2 d=9+16 d=√25 d=5

OpenStudy (agent0smith):

Correct, and that d is displacement. Now use velocity = displacement/time

OpenStudy (anonymous):

V= 5/10 v=-.5 m/s Thank you!!!!

OpenStudy (agent0smith):

Welcome :)